OpenStudy (hellomynameiskam):

Which pair of ratios is proportional? A.6/21 and 2/42 B.25/100 and 50/150 C. 18/25 and 6/12 D. 7/14 and 20/40

OpenStudy (igreen):

Easiest way is to divide them and see if you get the same thing.

OpenStudy (igreen):

6 / 21 = ? 2 / 42 = ? 25 / 100 = ? 50 / 150 = ? 18 / 25 = ? 6 / 12 = ? 7 / 14 = ? 20 / 40 = ?

OpenStudy (igreen):

Whichever two give you the same is your answer..

OpenStudy (hellomynameiskam):

What if i get a decimal for the answer?

OpenStudy (hellomynameiskam):

@iGreen

OpenStudy (igreen):

You'll always get a decimal..just see if the decimals are the same.

OpenStudy (hellomynameiskam):

o

OpenStudy (hellomynameiskam):

ok*

OpenStudy (hellomynameiskam):

I got D, is that correct? @iGreen

OpenStudy (igreen):

Yep! They both equal 0.5

OpenStudy (hellomynameiskam):

Okay. Thanks

OpenStudy (igreen):

No problem.

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'.